OpenStudy (anonymous):

5(w-3)=-15. using distributive i got w=0

OpenStudy (anonymous):

foid 5w-15=-15 5w= -15+15 w =0/5 w =0

OpenStudy (anonymous):

u have correct answer

OpenStudy (anonymous):

yay thanks

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

yw